### v5

@@ 4.1 Ascribing Intent to Other Objects
Earlier, we arrived at the following graph in which Bob’s intervention was given by $int$.
$int$$obs$$rain$
What if a third person Larry puts the coat on Alice? Bob may observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$. To account for this, Bob can construct a causal graph as below (with $b.$ representing Bob and $l.$ representing Larry).
$b.\ int$$l.\ int$$obs$$rain$
Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself. However, now we can see that Bob must also construct an identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$. For an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [56], to satisfy the incentive precondition for causal identity.
@@ 4.2 Preconditions
There are preconditions for the existence of a causal identity. First, the vocabulary $\mathfrak{v}_{\mathfrak{a}}$ of an organism $\mathfrak{a}$ must be of sufficient scale to ensure that observations are distinguishable from interventions.
Second, there must be an incentive to construct it. Inference is only possible if some states are preferable to others. One cannot derive what “ought” to be from a statement of what “is”. Natural selection provides a notion of what ought to be, by eliminating anything which ought not.
@@ 4.3 Realising Lower Order States And Higher Order Meta Representations
Each policy an organism $\mathfrak{o}$ learns implies $\mathfrak{v}$-tasks. A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [68, 56] of sign, referent and interpretant. We formalise this as a “protosymbol” [^11] Proto because it is something more primitive than a symbol as conceived of by Peirce. system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$ (see definitions 6 and 7).
Tasks exist in a “generational hierarchy”. They are not mutually exclusive. Higher level tasks are more general, and have fewer policies because only very weak policies could complete them. Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations. Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [69]. Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
@@ 5 Multi-Layered Self-Organisation
As the vocabulary and capacity for w-maxing scale, a greater variety of concepts can be learned. Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [39]. This lets us frame the construction of embodied selves in developmental [70] and evolutionary terms.
@@ 5.1 The First Order Self
A first order self (1ST henceforth, formalised in definition 13) allows an organism to discern the consequences of its actions. This serves as the locus of self related information processing and experience [39], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction. 1ST order selves amount to reafference [71, 72], observed in mammals and insects.
Figure 1: Visual intuition for a 1ST order self. The organism constructs a causal identity for itself and can relate that to observed events.
A 1ST order self may require centralisation and a “solid brain” with a persistent structure [73] (an individual human), as opposed to a “liquid brain” (e.g. an ant colony or a population of humans).
@@ 5.2 The Second Order Selves
Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes [56]). This is called a second order self (2ND henceforth). We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense [74, 75]. Grice argued that communication is about the inference of intent. If person $\mathfrak{a}$ and $\mathfrak{b}$ are talking, then the meaning $m_{\mathfrak{a}}$ of what $\mathfrak{a}$ says is whatever $\mathfrak{a}$ intends $\mathfrak{b}$ understand. The meaning $m_{\mathfrak{b}}$ that $\mathfrak{b}$ understands is whatever $\mathfrak{b}$ thinks $\mathfrak{a}$ wants $\mathfrak{b}$ to think. $\mathfrak{b}$ has understood what $\mathfrak{a}$ means if $m_{\mathfrak{b}}$ approximates $m_{\mathfrak{a}}$. This can happen only if $\mathfrak{a}$ can predict with reasonable accuracy what $\mathfrak{b}$ thinks $\mathfrak{a}$ thinks, and $\mathfrak{b}$ can predict what $\mathfrak{a}$ thinks $\mathfrak{b}$ will think upon hearing an utterance. In other words, both $\mathfrak{a}$ and $\mathfrak{b}$ must have 2ND order selves that are good approximations. Yes, there are other aspects to communication.
@@
Where a 1ST order self might allow one to observe a cat and form plans regarding causal interactions with the cat, a 2ND order self would allow one to be consciously aware of the cat for the purpose of reasoning and report. One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^13] This is ‘meta-self-reflexive consciousness’ as some have described it [76]..
Figure 2: Visual intuition for a 2ND order self. The organism constructs a causal identity for itself, and for another object, and the causal identity for the other object includes a prediction of the organism itself from that object’s perspective. This would occur, for example, if one was a predator trying to predict the movements of prey in response to one’s own actions. It is an extension of the 1ST order self.
More formally using the notation given in the appendix [40] (see quick reference guide in section 2), assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour. Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part. Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part. What is important here is that each organism’s intent is to some extent inferred by the other, changing the sorts of policies that are fit. For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [77] [^14] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [56], and transient relations, information asymmetry and other factors can make deceit a more attractive option.. Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [56]. To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent. The incentive precondition explains why $\mathfrak{a}$ would form such intent (co-operation is often advantageous), while how may be understood as follows:
@@
$\mathfrak{a}$ can use $c^{\mathfrak{ba}}_{\mathfrak{a}}$ to infer behaviour to which $\mathfrak{b}$ will ascribe the intent to communicate $m_{\mathfrak{a}}$, and $c^{\mathfrak{ab}}_{\mathfrak{b}}$ lets $\mathfrak{b}$ infer that this is what $\mathfrak{a}$ intends. The “utterance” Grice refers to is how $\mathfrak{a}$ affects $\mathfrak{b}$ in accord with earlier definitions. Put another way, $\mathfrak{a}$ encodes $m_{\mathfrak{a}}$ into its behaviour in a manner that $\mathfrak{b}$ can decode (their respective second order selves act as encoders and decoders). By encode and decode, we mean a loose approximation of $m_{\mathfrak{a}}$ is communicated. There are of course shortcuts, for example of $\mathfrak{a}$ and $\mathfrak{b}$ are of the same species then they likely have similar motives and experiences, and so the efficient thing for each to do would be to use its own intent as an approximation of what the other might think. However, that does not obviate the need for second order selves, it just makes them easier to realise.
@@ 5.3 The Third Order Selves
We can scale preconditions indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [76]. This is the awareness that one is self-aware. If self-awareness stems from 2ND order selves, then it follows that meta self-awareness requires 3RD. Formally a 3RD order self for $\mathfrak{a}$ reflecting off $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{b}$ is $c^{\mathfrak{baba}}_{\mathfrak{a}}$. It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
Figure 3: Visual intuition for a 3RD order self. The organism constructs a 2ND order self for its 2ND order self, and so becomes aware that it is aware.

### v6

@@ 4.1 Ascribing Intent to Other Objects
Earlier, we arrived at the following graph in which Bob’s intervention was given by intint.
$\mathrm{int}$$\mathrm{obs}$$\mathrm{rain}$ What if a third person Larry puts the coat on Alice? Bob may observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$. To account for this, Bob can construct a causal graph as below (with $b.$ representing Bob and $l.$ representing Larry).
$b.\,\mathrm{int}$$l.\,\mathrm{int}$$\mathrm{obs}$$\mathrm{rain}$ Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself. However, now we can see that Bob must also construct an identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$. For an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ Bennett (2023b), to satisfy the incentive precondition for causal identity.
@@
A likelihood ratio bound shows that decoder mismatch can hide minds. If every mind-like hypothesis in an evaluator’s model fits an observed trace worse than a noise hypothesis by a large factor, Bayes rule forces the posterior away from mind (see Box 3 and ESM for technical details).
@@ 4.2 Preconditions
There are preconditions for the existence of a causal identity. First, the vocabulary $\mathfrak{v}_{\mathfrak{a}}$ of an organism $\mathfrak{a}$ must be of sufficient scale to ensure that observations are distinguishable from interventions. Second, there must be an incentive to construct it. Inference is only possible if some states are preferable to others. One cannot derive what “ought” to be from a statement of what “is”. Natural selection provides a notion of what ought to be, by eliminating anything which ought not. In a creature that can die, death grounds meaning Bennett (2025a).
@@ 4.3 Realising Lower Order States And Higher Order Meta Representations
Each policy an organism $\mathfrak{o}$ learns implies $\mathfrak{v}$-tasks. A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies. This resembles Peircean semiosis Atkin (2024); Bennett (2023b) of sign, referent and interpretant. We call the resulting family of reusable tasks a “protosymbol” [^15] Proto because it is more primitive than a Peircean symbol. system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$. Intuitively, a protosymbol is a behavioural template that can be reused across contexts (full definition in ESM). Tasks exist in a “generational hierarchy”. They are not mutually exclusive. Higher level tasks are more general, and have fewer policies because only very weak policies could complete them. Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations. Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations Goertzel (2006). Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
Box 2: Theorem 3. The Psychophysical Principle of Causality. One can learn causality by learning the objects and properties that fit a causal relation, or one can learn causal relations that fit a set of presupposed objects and properties. However if you want generalisation optimal learning, you cannot hard code a quality neutral ontology (no presupposed objects and properties). Only the good or bad signal can be a fixed representational primitive from which more abstract classifications are constructed as needed. This is because valence is the signal that determines how representations are judged, and so any fixed abstraction which is not a direct consequence of that signal only serves to obfuscate the signal and impose a floor on variational free energy Bennett (2024a). Continued existence and its lack determine what is attractive to whatever remains. Hence causal relation of valence is given. In a living system a causal identity is constructed of valence. It is a pattern in valence. A sufficiently complex mixture of repulsive and attractive signals, can together amount to a qualitatively neutral classification like “the colour red”. Attraction and repulsion is how an organism carves up its world into objects and properties. Objects and properties are made of valence, and so are qualitative rather than quantitative information. This is the informal version of The Psychophysical Principle of Causality Bennett (2025a). Here we reformulate this idea as an equivalent mathematical theorem, and prove (rather than just argue) that optimality demands valence first learning (at least, within the formal confines of the Stack Theory). We number this as Theorem 3 to match the numbering in the Supplementary Information. Theorem 3 is a corollary to The Law of the Stack (see ESM), but is itself a novel contribution of this paper. Definition (hard primitive commitments). Let $\alpha$ be a task and let $\Pi_{\alpha}$ be its set of correct policies. Let $C\subseteq L_{\mathfrak{v}}$ be a nonempty set of commitments. The restricted policy class that bakes in $C$ is $\Pi_{\alpha}^{C}:=\{\pi\in\Pi_{\alpha}:\forall c\in C,c\subseteq\pi\}.$ We call $C$ a set of hard primitives when we enforce this restriction regardless of the evidence. Setup. Let $\mu$ be a viability statement encoding the most basic valence partition. Viable versus not viable. For any policy $\pi$, define the viable continuation set $\Omega_{\pi}:=E_{\pi}\cap E_{\mu}.$ This is the set of viable future completions still compatible with $\pi$. Define the free energy proxy in bits $F_{2}(\pi):=\log_{2}|E_{\mu}|-\log_{2}|\Omega_{\pi}|.$ This equals $D_{\mathrm{KL},2}(q_{\pi}\|p_{\mu})$. Here $q_{\pi}$ is uniform on $\Omega_{\pi}$ and $p_{\mu}$ is uniform on $E_{\mu}$. Theorem 3 (The Psychophysical Principle of Causality). Assume $C\neq\emptyset$ contains at least one commitment that is not forced by viability. This means there exists $o\in E_{\mu}$ such that $c\not\subseteq o$ for some $c\in C$. Then any policy that is forced to include $C$ has strictly higher $F_{2}$ whenever $C$ excludes at least one viable continuation that would otherwise remain possible. So a generalisation optimal learner cannot keep such non valence primitives fixed. Proof sketch. Adding a commitment intersects the viable continuation set with a smaller extension. That can only shrink $\Omega_{\pi}$. Since $F_{2}(\pi)$ is monotone in $|\Omega_{\pi}|$, the free energy increases strictly if any viable continuation is removed. See ESM Theorem 3 for the full proof. Interpretation. The only safe primitive is the qualitative viability signal. That is what we call valence. Quality neutral representations must therefore be constructed later, as learned causes of valence.
@@ 5 Multi-Layered Self-Organisation
As the vocabulary and capacity for w-maxing scale, a greater variety of concepts can be learned. Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed Bennett (2023a). This lets us frame the construction of embodied selves in developmental Ciaunica and Crucianelli (2019) and evolutionary terms.
@@ 5.1 The First Order Self
A first order self (1ST henceforth) allows an organism to discern the consequences of its actions. In our formalism it is the weakest causal identity that distinguishes self generated interventions from matched observations. The full definition and proof are in the ESM. This serves as the locus of self related information processing and experience Bennett (2023a), allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction. 1ST order selves amount to reafference Merker (2007); Barron and Klein (2016), observed in mammals and insects.
A formal role for the 1ST order self.
If an organism can improve expected utility by treating intervention episodes differently from observation episodes, then some internal variable must encode that difference. Otherwise the organism would be forced to map the same embodied statement to the same internal state, and so could not act differently in the two cases.
Under the scale and incentive preconditions already described, the candidate set of self intervention signatures is nonempty. A w-max learning rule then selects a maximally weak signature from that set. Any w-maximiser is inclusion minimal among candidates. If a strictly smaller signature existed, it would have a strictly larger extension and would have been preferred. This is the operational content of reafference in our framework.
Why self is a fixed point of w-maxing.
A short proposition shows that if there exists a minimal self intervention tag that is implied by every viable policy, and that tag is itself viable, then it is the weakest viable policy and w-maxing converges to it. See Box 3 for the sketch and see the ESM proposition titled “Self fixed point” for the full proof.
A 1ST order self may require centralisation and a “solid brain” with a persistent structure Solé et al. (2019) (an individual human), as opposed to a “liquid brain” (e.g. an ant colony or a population of humans). A solid brain would be required for the signals that contribute to a first order self to be co-instantiated in proper time Bennett (2026a), and causally affect one another Bennett (2026c).
@@ 5.2 The Second Order Selves
Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes Bennett (2023b)). This is called a second order self (2ND henceforth). We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense Grice (1957, 1969). Grice argued that communication is about the inference of intent. If person $\mathfrak{a}$ and $\mathfrak{b}$ are talking, then the meaning $m_{\mathfrak{a}}$ of what $\mathfrak{a}$ says is whatever $\mathfrak{a}$ intends $\mathfrak{b}$ understand. The meaning $m_{\mathfrak{b}}$ that $\mathfrak{b}$ understands is whatever $\mathfrak{b}$ thinks $\mathfrak{a}$ wants $\mathfrak{b}$ to think. $\mathfrak{b}$ has understood what $\mathfrak{a}$ means if $m_{\mathfrak{b}}$ approximates $m_{\mathfrak{a}}$. This can happen only if $\mathfrak{a}$ can predict with reasonable accuracy what $\mathfrak{b}$ thinks $\mathfrak{a}$ thinks, and $\mathfrak{b}$ can predict what $\mathfrak{a}$ thinks $\mathfrak{b}$ will think upon hearing an utterance. In other words, both $\mathfrak{a}$ and $\mathfrak{b}$ must have 2ND order selves that are good approximations. Yes, there are other aspects to communication.
@@
Where a 1ST order self might allow one to observe a cat and form plans regarding causal interactions with the cat, a 2ND order self would allow one to be consciously aware of the cat for the purpose of reasoning and report. One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^17] This is ‘meta-self-reflexive consciousness’ as some have described it Morin (2006)..
More formally using the notation given in the appendix Bennett (2025a) (see quick reference guide in section 2), assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour. Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part. Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part. What is important here is that each organism’s intent is to some extent inferred by the other, changing the sorts of policies that are fit. For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate Alexander et al. (2022) [^18] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others Bennett (2023b), and transient relations, information asymmetry and other factors can make deceit a more attractive option.. Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) Bennett (2023b). To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent. The incentive precondition explains why $\mathfrak{a}$ would form such intent (co-operation is often advantageous), while how may be understood as follows:
@@
$\mathfrak{a}$ can use $c^{\mathfrak{ba}}_{\mathfrak{a}}$ to infer behaviour to which $\mathfrak{b}$ will ascribe the intent to communicate $m_{\mathfrak{a}}$, and $c^{\mathfrak{ab}}_{\mathfrak{b}}$ lets $\mathfrak{b}$ infer that this is what $\mathfrak{a}$ intends. The “utterance” Grice refers to is how $\mathfrak{a}$ affects $\mathfrak{b}$ in accord with earlier definitions. Put another way, $\mathfrak{a}$ encodes $m_{\mathfrak{a}}$ into its behaviour in a manner that $\mathfrak{b}$ can decode (their respective second order selves act as encoders and decoders). By encode and decode, we mean a loose approximation of $m_{\mathfrak{a}}$ is communicated. There are of course shortcuts, for example of $\mathfrak{a}$ and $\mathfrak{b}$ are of the same species then they likely have similar motives and experiences, and so the efficient thing for each to do would be to use its own intent as an approximation of what the other might think. However, that does not obviate the need for second order selves, it just makes them easier to realise. Formal and experimental results supporting our claims regarding 2ND order selves are given in the ESM.
A minimal separation task.
Consider a speaker who must transmit a private bit to a listener. The listener can be in one of two decoder modes. One decodes literally and the other flips the bit. If the speaker cannot condition on any evidence about which decoder is active, its success is at most one half. If the speaker can send one probe signal and observe the listener’s response, it can infer the decoder and succeed with probability one. That conditioning variable is the smallest possible second order self. It is an internal handle for how the audience is interpreting the speaker.
Decoder mismatch in Monte Carlo benchmarks.
Monte Carlo benchmarks in ESM Experiment S2 show that a couple of short probes largely eliminate the cost of decoder mismatch. Passive signalling stays near chance, while probing approaches oracle performance. See Box 4 for the headline results, and see ESM Experiment S2 for the full protocol and plots.
@@ 5.3 The Third Order Selves
We can scale preconditions indefinitely. For the purposes of this paper we stop at 3RD order because this is where a self can extend across time as a stable narrative. It is where an agent can treat its present actions as evidence about its future behaviour. It is also where the agent can deliberately bind itself to reduce what it could do later. This is what makes long horizon cooperation and trust possible.
Formally, a 3RD order self for $\mathfrak{a}$ relative to an audience $\mathfrak{b}$ is the iterated causal identity c^baba_a. It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$. $\mathfrak{a}$’s action now changes what $\mathfrak{b}$ expects of $\mathfrak{a}$ later, and $\mathfrak{a}$ knows that $\mathfrak{b}$ knows that $\mathfrak{a}$ knows.
A formal separation in the ESM shows why narrative, 3RD order selves matter for trust. Even with perfect audience modelling, nonbinding promises do not support trust in equilibrium when exploitation remains available. If an agent can publicly bind itself by paying a one time cost that removes an exploit option, a trust equilibrium exists. Monte Carlo benchmarks in ESM Experiment S3 show the same pattern across randomised trust games. See Box 3 for the formal sketch and Box 4 for the experimental headline results. See ESM Experiment S3 for the full protocol and plots.
Box 3: Other formal proofs in support of the main text claims The main text uses proof results besides Theorem 3. They constrain and support the narrative arguments of this paper. Here we give only the sketch level version. Complete proofs are in the ESM. Reminder. Weakness maximisation, called w-maxing, means selecting a correct policy with maximal weakness $w(\pi)=|E_{\pi}|$. In plain English, it keeps the largest set of future refinements open while staying correct. To tie this to related biological frameworks on cognition spaces Solé et al. (2026), weakness corresponds to the size of the “cognitive light cone” of an embodied system. Result 1. w-maxing learns causal intervention tags. Given matched intervention and observation episodes, a w-maximised causal identity exists in any finite embodied language and is inclusion minimal among candidates. In plain English, it is the least committal signature that still separates self caused change from external change. This proves an earlier but less formal claim Bennett (2023a). Result 2. Decoder mismatch can hide minds. A Bayesian likelihood ratio bound shows that an evaluator can be forced toward a noise hypothesis when every mind-like hypothesis in its model class fits the observed trace worse than noise by a constant factor. In other words, if you decode the signal with the wrong key, Bayes rule will tell you there is no sender [^19] This creates mind-blindness and relates to The Fermi Paradox Bennett (2022a).. Result 3. The 1ST order self is a w-max fixed point. If there exists a minimal self intervention tag that is implied by every viable policy and is itself viable, then it is the weakest viable policy and a w-max learner converges to it. In plain English, reafference becomes a stable self tag when survival depends on it. It is a mathematical result that predicts the reafference we observe in living organisms. Result 4. Gricean report and trust force higher order selves. In Gricean communication, what counts as a correct report depends on the listener’s inference, not a fixed signal to world mapping. We show that any signalling policy that is robust across variation in an audience’s epistemic state must condition on an audience model. In our formalism, that audience model is a 2ND order self. We also give a minimal commitment game in which even perfect audience modelling cannot sustain trust when exploitation remains available. If an agent can publicly bind itself by paying a one time cost that removes an exploit option, a trust equilibrium exists. In other words, reportable access is not just the ability to say a thing. It is saying the thing knowing a listener will understand, by predicting what they predict you intend by what you say, and sometimes even binding yourself so they can believe you (because they will predict your self bind and observe your behaviour is consistent with it). Taken together, Results 3 and 4 make the self hierarchy a derived necessity. It is the minimal structure that survives under generalisation optimal adaptation when the niche includes self intervention, report, and trust. See the ESM propositions on causal identity minimality, decoder mismatch, self fixed points, Gricean report, and credible commitment.